OpenStudy (anonymous):

the formula y = 2.205x converts a mass x in kilograms to a weight y in pounds. A sports car has a mass of 1270 kilograms. what is its weight in pounds?

OpenStudy (anonymous):

x=y/2.205 x=1270/2.205

OpenStudy (anonymous):

y=1270/2.205

OpenStudy (anonymous):

oh sry...first one was correct..i think

OpenStudy (anonymous):

22.05*1270

OpenStudy (anonymous):

2.205*1270

OpenStudy (anonymous):

thank you guys :DD i love you guys :DD

OpenStudy (anonymous):

thank you guys :DD i love you guys :DD